# The Enigma of Li-Rich Giants and its Relation with Temporal Variations Observed in Radial Velocity and Stellar Activity Signals

### v1

@@ Abstract
Context. Despite the large number of studies focused on the characterization of Li-rich stars and understanding of the mechanisms leading to such enrichment, their origin remains a mystery.
Aims. Magnetic activity, in particular the phenomena usually associated with it, such as spots and plages, and the Li abundance (A(Li)) of stars, are in general thought to be connected. However, as of today it is unclear just how. In this work, we study a sample of young but evolved intermediate mass red giants, inhabitants of open clusters where planets have been searched. We aim at using radial velocity (RV) and stellar activity indicator signals to look for relations between Li abundances and stellar activity/variability.
Methods. We explore how the standard deviation (STD), peak to peak amplitude (PTP), mean and median of typical stellar activity indicators (BIS, FWHM, , and H index) change as a function of the Li content of 82 red giants. Furthermore, we compute weighted Pearson Correlation Coefficients () between time series of RV measurements and the stellar activity indicators for the stars in our sample. To aid our results, we also study Generalized Lomb-Scargle Periodograms (GLSP) to capture possible significant periodic temporal variations in our data.
Results. Our analysis indicates that the STD and PTP of BIS and FWHM, the mean and median of the H index and increase exponentially with A(Li) in our sample of red giants. Significant temporal variations and correlations between RVs and activity indicators also tend to be found preferentially for stars where high A(Li) is observed. Most of the Li-rich stars in our sample either show strong correlations of RV with at least one of the stellar activity indicators or reveal significant periodic temporal variations in their stellar activity indicators GLSPs consistent with those found for RV.
@@
Low () to intermediate mass stars () (Iben 1967a, b), before entering the red giant branch (RGB) of their evolution, go through a mixing episode called the first dredge up (FDU), which destroys most, if not all, of their surface lithium (Li). Stars as massive as the Sun, according to standard stellar evolution models, should not exhibit a surface lithium abundance, A(Li)111, exceeding 1.5 dex after the FDU. In general, observations concur this but 1% of observed giants (e.g Brown et al. 1989; Kumar et al. 2011; Liu et al. 2014; Casey et al. 2016; Kirby et al. 2016; Deepak & Reddy 2019; Gao et al. 2019; Pabst et al. 2020; Martell et al. 2021) present abnormally high A(Li) with the highest abundance observed to date in a red giant star being 5.62 dex (Kowkabany et al. 2022). These stars are known as Li-rich giants.
@@
It is also generally accepted that higher than normal A(Li) is common among magnetically active stars, even if the reason for this is still poorly understood. This has been evidenced for different types of stars namely dwarfs (Favata et al. 1996; Zboril et al. 1997), magnetically peculiar stars (see Lyubimkov 2016 for a review), active binaries and pre-main-sequence stars (Pallavicini et al. 1992), active main sequence and zero-age main sequence stars (Xing et al. 2021) and active giant stars (Fekel & Balachandran 1994; Sneden et al. 2022).
In this work, we are interested in studying the possible connection between Li enhancement in red giant stars and their stellar activity. For this we start by looking for relations between several descriptive statistics of stellar activity indicators and A(Li) in a sample of red giant stars, some of which are lithium rich, where planets have been searched. These stars are inhabitants of open clusters, which is particularly favourable, as we are able to better constrain their evolutionary status. To strengthen our analysis we also look for correlations as well as periodic temporal variations in stellar activity signals that may connect variability observed in radial velocity (RV) to stellar activity. The present article is organized as follows, in Sect. 2 we provide an overview of our data. In Sect. 3 we explain the methods incurred in the data analysis. In Sect. 4 we discuss our results, and the main conclusions are laid out in Sect. 5.
@@ 2 Observations and Sample
Our sample is comprised of young but evolved intermediate mass red giants resident in 10 different open clusters (IC4651, IC4756, NGC2345, NGC2423, NGC3114, NGC3532, NGC3680, NGC4349, NGC5822 and NGC6705). These stars have been observed over the years with the HARPS spectograph (Mayor et al. 2003) at the ESO 3.6-metre telescope (R=115 000) and the CORALIE spectrograph (Queloz et al. 2000; Udry et al. 2000) at the Swiss 1.2-metre Leonhard Euler Telescope at ESO’s La Silla Observatory (R=60 000), as part of a program to search for exoplanets. For a complete description of the RV surveys see Lovis & Mayor (2007); Delgado Mena et al. (2018); Delgado Mena et al. (2023). The stellar parameters and Li abundances for this sample are presented in Tsantaki et al. (2023), obtained using the same spectra employed here.
@@
In this work, the data available from CORALIE will not be treated. CORALIE data carries a much higher associated uncertainty in the measurements than HARPS data, due to its significantly lower resolution (R). Furthermore, CORALIE data was only collected for a subset of the stars in our sample, and the measurements for each star are few in number. For these reasons it is preferable to use only HARPS data.
@@
Only stars with 5 or more observational data points were considered for analysis. This is such to increase the credibility of our results, while including the highest amount of observed stars possible. In total, 82 stars were analysed out of which 15 are canonically Li-rich, that is, show A(Li)1.5 dex. If a Population I star forming with a metallicity alike that of the Sun has in the beginning of its life A(Li)3.3 dex, that is, the approximate A(Li) of the interstellar medium, accounting for the dilution occurring during the FDU, in the red giant phase of its evolution it should be left with a maximum of A(Li)1.5 dex (e.g. Charbonnel & Balachandran 2000). Studying stars in open clusters, however, allows for a more precise estimate of this limit. Open clusters are ideal laboratories for studying Li enhancement given that they provide natural constraints on mass, composition, age and evolutionary state of their members. With such information it is possible to estimate the amount of Li that all stars in a given cluster should have at their current age using stellar evolution models. This allows us to better constrain what makes a star in a given cluster “Li-rich” (see Sun et al. 2022). Tsantaki et al. (2023) conducted such a study, providing estimations for the maximum A(Li) that stars in 32 different open clusters should have in their current evolutionary stage. Stars will be considered Li-rich according to the limits defined in Table 1 and the canonical definition. Special attention will also be given to stars that show high Li content in comparison with inhabitants of the same cluster in the same evolutionary stage, these will be deemed ‘Li-Rich-in-Context’ in accordance with Tsantaki et al. (2023) and Sun et al. (2022). All relevant stars come presented in Table 2. Some of these stars have been previously studied in the literature. A thorough discussion concerning the possible nature of the Li enrichment observed in these stars has been provided in Tsantaki et al. (2023), and is summarized in Table 2.
@@
Notes: () This work; () Possible non-member: Katime Santrich et al. (2013). () As derived in Katime Santrich et al. (2013). () As derived in Smiljanic, R. et al. (2009). () As derived in Luck (1994). () Non-member: Mermilliod et al. (2008). () Possible non-member: Herzog et al. (1975); Frinchaboy & Majewski (2008); Bagdonas et al. (2018). () As derived in Holanda et al. (2022). () As derived in Holanda et al. (2019). () As derived in Peña Suárez et al. (2018).
@@ 3 Methods
To probe if the high A(Li) that we observe in red giant stars is connected to their stellar activity we explore two methods. First, we study how Li abundances relate to these five stellar activity indicators:
@@
In particular, we look for relations between A(Li) and the standard deviation (STD) and peak to peak amplitude (PTP) of all stellar activity indicators as well as the mean and median of H. Only statistically significant models at a 99% confidence level (with p-values of the F-test ) will be discussed.
@@
Secondly, we would like to investigate if a possible connection between A(Li) and stellar activity is manifested in RV variability. So, we seek correlations between stellar activity indicators and RV measurements (see Figure 5). These will be computed according to the Pearson Correlation Coefficient, , which is here weighted by the uncertainty in the measurements. P-values for each correlation will also be provided to attest the statistical significance of the correlations. A p-value bellow 0.05 suggests that the correlation is significantly different from 0 at a 95% confidence level. In addition, periodic temporal variations of the RV measurements of these stars and the stellar activity indicators will be analysed by drawing Generalized Lomb-Scargle Periodograms (GLSP) (Zechmeister & Kürster 2009), to evaluate if they connect to the presence of Li.
@@
When searching for planets it is important to devise methods to distinguish planetary signals from temporal variability that may arise from the phenomena appearing at the surface of the star as a consequence of stellar magnetic activity. The BIS and FWHM values of each observation are provided by the HARPS data reduction pipeline as part of the process to calculate the RV of a given spectrum. Correlations found between RV and respective BIS and FWHM measurements indicate that periodical variability noticed in RV signals may originate from stellar activity or stellar surface inhomogeneities, rather than planets. Variability observed in BIS and FWHM over time with approximately the same period as that observed in RV, can also further corroborate this. These types of studies have often been conducted to probe whether or not a planet really exists in orbit of a certain star (e.g. Queloz et al. 2001; Delgado Mena et al. 2018).
@@
In studying how stellar activity impacts an RV signal, it is common to analyse spectral lines that serve as an activity proxy due to their well known correlations with the existence of active regions in the surface of the star. One such spectral line is the Hα line. Here, we determine Hα index values from spectra using ACTIN 2222https://github.com/gomesdasilva/ACTIN2 (Gomes da Silva et al. 2018, 2021). This tool to calculate stellar activity indices provides two measurements for the Hα index for two different central bandpasses: 0.6 Å and 1.6 Å (see Figure 6). Since it has been mostly used for dwarf stars (Gomes da Silva et al. 2022), there is a lack of an extensive study for stars as the ones in our sample, and it is not clear which central bandpass is optimal for the study of these stars. Recently, in analysing the strong RV signals of some of the stars in this sample (Delgado Mena et al. 2023), we found that H16 seems to be a more reliable activity indicator for these evolved stars. But this conclusion was based on a small sample. As such, and since the two bandpasses probe the stellar atmosphere at different layers being sensitive to different phenomena, in our analysis both are used.
@@
We further extend our search for correlations and periodical variations to the Teff. Manifestations of magnetic activity, much like spots and plages are usually accompanied by temperature variations. Spots are cooler than their surrounding photosphere, while plages are hotter than their surrounding chromosphere. During periods of intense magnetic activity it is possible that, if the Teff is affected, the variations observed could correlate with the activity induced variations (e.g. Yana Galarza et al. 2019). We use the stellar spectral analysis package, FASMA333https://github.com/MariaTsantaki/FASMA-synthesis (Tsantaki et al. 2018), to determine the Teff from individual stellar spectra. FASMA relies on the radiative transfer code MOOG444https://www.as.utexas.edu/~chris/moog.html (Sneden 1973) to model the atmosphere and create a synthetic spectrum for a set of input parameters (surface gravity, ; iron metallicity, [Fe/H]; macro-turbulence velocity ; and projected rotational velocity ) of a given star. The synthetic spectrum is then statistically compared to the observed spectrum of the star and the parameters are iteratively adjusted using a non-linear least-squares algorithm to find the set that best fits observations. To capture reliable variations in Teff, when computing it, we fix all other input stellar parameters, and leave Teff to be computed freely for each individual stellar spectra. Stellar parameters for the stars in our sample have been derived in Tsantaki et al. (2023) also using FASMA.
@@ 4 Results and Discussion
As presented in Table 2 and in accordance with Mermilliod et al. (2008), the star NGC3532No649 is not a member of cluster NGC3532 making the calculation of its mass and evolutionary stage uncertain. For this reason, it will not be included in the following discussion despite it being Li-rich.
4.1 Is the Lithium Abundance of Red Giants related to Stellar Activity?
We start by analysing the dispersion (STD and PTP) of BIS, FWHM, T and H to probe potential connections between activity variability and A(Li). We also test if the mean and median of Hα, indicative of a star’s activity level, is related to A(Li). For the computation of these relations we only use stars with A(Li) as this is the limit of FASMA’s minimization algorithm in the determination of A(Li) (Tsantaki et al. 2023). Using these limit values would be a source of bias. We find that the STD and PTP of BIS and FWHM increase exponentially with A(Li) and the mean and median values of H06 also increase exponentially with A(Li). In particular, the vast majority of stars with high H06 index are Li enriched. On the other hand, we do not find A(Li) to be related to H16. These results suggest that the activity levels of the stars in our sample seem to increase with increasing A(Li). These results are in agreement with Sneden et al. (2022) who found that Li-rich giants are more chromospherically active when in comparison to their Li poor counterparts, as they are more likely to show strong He I10830 absorptions. We also find that correlates well with dispersion of BIS and FHWM as well as the mean and median of the H06 index. This is indicative that more magnetically active stars also tend to rotate faster. Gonçalves et al. (2020) also arrived at a similar conclusion. The best fit relations are listed in Table 3. Plots of the relations found for STD of BIS and FWHM as well as the mean of H06 are shown in Figure 1. The plots for PTP of BIS and FWHM and the median of H06 look similar to the ones shown in Figure 1 and are presented in Appendix B. Also in Appendix B are plots of the mean and median H16 as a function of A(Li), for which no clear relation was found. In addition to the results of the F-test, in Table 3 we also provide R2 values, with better fits having R2 closer to 1. R2 values for the relations we find are in general low as there is significant scatter around the lines of best fit.

### v2

@@ Abstract
Context. Despite the large number of studies focused on the characterisation of Li-rich stars and understanding the mechanisms leading to such enrichment, their origin remains a mystery.
Aims. Magnetic activity, particularly the phenomena usually associated with it (e.g. spots and plages), and the Li abundance (A(Li)) of stars, are in general thought to be connected. As of today, however, just how they are connected is unclear. In this work, we study a sample of young but evolved intermediate-mass red giants that are inhabitants of open clusters where planets have been searched for. Our aim is to use radial velocity (RV) and stellar activity indicator signals to look for relations between Li abundances and stellar activity or variability.
Methods. We explored how the standard deviation (STD), peak-to-peak amplitude (PTP), mean, and median of typical stellar activity indicators (BIS, FWHM, , and H index) change as a function of the Li content of 82 red giants. Furthermore, we computed weighted Pearson correlation coefficients () between time series of RV measurements and the stellar activity indicators for the stars in our sample. To aid our results, we also studied generalized Lomb-Scargle periodograms (GLSP) to capture possible significant periodic temporal variations in our data.
Results. Our analysis indicates that the STD and PTP of BIS and FWHM, the mean and median of the H index, and increase exponentially with A(Li) in our sample of red giants. Significant temporal variations and correlations between RVs and activity indicators also tend to be found preferentially for stars where high A(Li) is observed. Most of the Li-rich stars in our sample either show strong correlations of RV with at least one of the stellar activity indicators or reveal significant periodic temporal variations in their GLSPs of stellar activity indicators that are consistent with those found for RV.
@@
Before entering the red giant branch (RGB) of their evolution, low- () to intermediate-mass stars () (Iben 1967a, b) go through a mixing episode called the first dredge up (FDU), which destroys most, if not all, of their surface lithium (Li). Stars as massive as the Sun, according to standard stellar evolution models, should not exhibit a surface lithium abundance, A(Li)111, exceeding 1.5 dex after the FDU. In general, observations concur with this, but 1% of observed giants (e.g. Brown et al. 1989; Kumar et al. 2011; Liu et al. 2014; Casey et al. 2016; Kirby et al. 2016; Deepak & Reddy 2019; Gao et al. 2019; Pabst et al. 2020; Martell et al. 2021) present abnormally high A(Li), with the highest abundance observed to date in a red giant star being 5.62 dex (Kowkabany et al. 2022). These stars are known as Li-rich giants.
@@
It is also generally accepted that higher than normal A(Li) is common among magnetically active stars, even if the reason for this is still poorly understood. This has been shown in different types of stars, including dwarfs (Favata et al. 1996; Zboril et al. 1997); magnetically peculiar stars (see Lyubimkov 2016 for a review); active binaries and pre-main-sequence stars (Pallavicini et al. 1992); active main-sequence and zero-age main-sequence stars (Xing et al. 2021); and active giant stars (Fekel & Balachandran 1994; Sneden et al. 2022).
In this work, we are interested in studying the possible connection between Li enhancement in red giant stars and their stellar activity. As such, we started by looking for relations between several descriptive statistics of stellar activity indicators and A(Li) in a sample of red giant stars, some of which are lithium rich, where planets have been searched. These stars are inhabitants of open clusters, which is particularly favourable, as we are able to better constrain their evolutionary status. To strengthen our analysis, we also looked for correlations as well as periodic temporal variations in stellar activity signals that may connect the variability observed in radial velocity (RV) to stellar activity. The present article is organised as follows: In Sect. 2, we provide an overview of our data. In Sect. 3, we explain the methods incurred in the data analysis. In Sect. 4, we discuss our results, and the main conclusions are laid out in Sect. 5.
@@ 2 Observations and sample
Our sample is composed of young but evolved intermediate-mass red giants residing in ten different open clusters (IC4651, IC4756, NGC2345, NGC2423, NGC3114, NGC3532, NGC3680, NGC4349, NGC5822, and NGC6705). These stars have been observed over the years with the HARPS spectrograph (Mayor et al. 2003) at the ESO 3.6-metre telescope (R=115 000) and the CORALIE spectrograph (Queloz et al. 2000; Udry et al. 2000) at the Swiss 1.2-metre Leonhard Euler Telescope at ESO’s La Silla Observatory (R=60 000) as part of a programme to search for exoplanets. (For a complete description of the RV surveys see Lovis & Mayor (2007); Delgado Mena et al. (2018); Delgado Mena et al. (2023).) The stellar parameters and Li abundances for this sample are presented in Tsantaki et al. (2023), who used the same spectra employed here.
@@
In this work, the data available from CORALIE are not treated. CORALIE data carry a much higher associated uncertainty in the measurements than HARPS data due to a significantly lower resolution (R). Furthermore, CORALIE data were only collected for a subset of stars in our sample, and the measurements for each star are few in number. For these reasons we preferred to use only HARPS data.
@@
Only stars with five or more observational data points were considered for analysis. This limitation increases the credibility of our results while including the most observed stars as possible. In total, 82 stars were analysed out of which 15 are canonically Li-rich, that is they show A(Li)1.5 dex. If a Population I star forming with a metallicity similar to that of the Sun has in the beginning of its life A(Li)3.3 dex, that is, the approximate A(Li) of the interstellar medium, accounting for the dilution occurring during the FDU, it should be left with a maximum of A(Li)1.5 dex in the red giant phase of its evolution (e.g. Charbonnel & Balachandran 2000). Studying stars in open clusters, however, allows for a more precise estimate of this limit. Open clusters are ideal laboratories for studying Li enhancement given that they provide natural constraints on the mass, composition, age, and evolutionary state of their members. With such information, it is possible to estimate the amount of Li that all stars in a given cluster should have at their current age using stellar evolution models. This allows us to better constrain what makes a star in a given cluster ‘Li-rich’ (see Sun et al. 2022). Tsantaki et al. (2023) conducted such a study, providing estimations for the maximum A(Li) that stars in 32 different open clusters should have in their current evolutionary stage. Stars are considered Li-rich according to the limits defined in Table 2 and the canonical definition. Special attention is given to stars that show high Li content in comparison to inhabitants of the same cluster in the same evolutionary stage. We deem such stars ‘Li-Rich-in-Context’, in accordance with Tsantaki et al. (2023) and Sun et al. (2022). All relevant stars are presented in Table 3. Some of these stars have been studied previously in the literature. A thorough discussion concerning the possible nature of the Li enrichment observed in these stars has been provided in Tsantaki et al. (2023), and it is summarised in Table 3.
@@
() This work; () Possible non-member: Katime Santrich et al. (2013). () As derived in Katime Santrich et al. (2013). () As derived in Smiljanic, R. et al. (2009). () As derived in Luck (1994). () Non-member: Mermilliod et al. (2008). () Possible non-member: Herzog et al. (1975); Frinchaboy & Majewski (2008); Bagdonas et al. (2018). () As derived in Holanda et al. (2022). () As derived in Holanda et al. (2019). () As derived in Peña Suárez et al. (2018).
@@ 3 Methods
To probe if the high A(Li) that we observe in red giant stars is connected to their stellar activity, we explored two methods. First, we studied how Li abundances relate to the following stellar activity indicators:
@@
In particular, we looked for relations between A(Li) and the standard deviation (STD) and peak-to-peak amplitude (PTP) of all stellar activity indicators as well as the mean and median of H. Only statistically significant models at a 99% confidence level (with p-values of the F-test ) are discussed.
@@
Secondly, we investigated if a possible connection between A(Li) and stellar activity manifests in RV variability. So, we sought correlations between stellar activity indicators and RV measurements (see Figure 5). These were computed according to the Pearson correlation coefficient, , which is weighted here by the uncertainty in the measurements. The p-values for each correlation are also provided to attest to the statistical significance of the correlations. A p-value bellow 0.05 suggests that the correlation is significantly different from zero at a 95% confidence level. In addition, periodic temporal variations of the RV measurements of these stars and the stellar activity indicators were analysed by drawing generalised Lomb-Scargle Periodograms (GLSPs (Zechmeister & Kürster 2009)) in order to evaluate if they connect to the presence of Li.
@@
When searching for planets, it is important to devise methods to distinguish planetary signals from temporal variability that may arise from the phenomena appearing at the surface of a star as a consequence of stellar magnetic activity. The BIS and FWHM values of each observation are provided by the HARPS data reduction pipeline as part of the process to calculate the RV of a given spectrum. Correlations found between RV and respective BIS and FWHM measurements indicate that periodical variability noticed in RV signals may originate from stellar activity or stellar surface inhomogeneities rather than planets. Variability observed in BIS and FWHM over time with approximately the same period as that observed in RV can also further corroborate this. These types of studies have often been conducted to probe whether or not a planet really exists in the orbit of a certain star (e.g. Queloz et al. 2001; Delgado Mena et al. 2018).
@@
In studying how stellar activity impacts an RV signal, it is common to analyse spectral lines that serve as an activity proxy due to their well-known correlations with the existence of active regions on the surface of a star. One such spectral line is the Hα line. Here, we determined Hα index values from spectra using ACTIN 2444https://github.com/gomesdasilva/ACTIN2 (Gomes da Silva et al. 2018, 2021). This tool to calculate stellar activity indices provides two measurements for the Hα index for two different central bandpasses: 0.6 Å and 1.6 Å (see Figure 6). Since it has been mostly used for dwarf stars (Gomes da Silva et al. 2022), there is a lack of an extensive study for stars such as the ones in our sample, and it is not clear which central bandpass is optimal for the study of these stars. Recently, in analysing the strong RV signals of some of the stars in this sample (Delgado Mena et al. 2023), we found that H16 seems to be a more reliable activity indicator for these evolved stars. But this conclusion was based on a small sample. As such, and since the two bandpasses probe the stellar atmosphere at different layers that are sensitive to different phenomena, we used both bandpasses in our analysis.
@@
We further extended our search for correlations and periodical variations to the Teff. Manifestations of magnetic activity, much like spots and plages are usually accompanied by temperature variations. Spots are cooler than their surrounding photosphere, while plages are hotter than their surrounding chromosphere. During periods of intense magnetic activity, it is possible that if the Teff is affected, the variations observed could correlate with the activity-induced variations (e.g. Yana Galarza et al. 2019). We used the stellar spectral analysis package FASMA555https://github.com/MariaTsantaki/FASMA-synthesis (Tsantaki et al. 2018) to determine the Teff from individual stellar spectra. FASMA relies on the radiative transfer code MOOG666https://www.as.utexas.edu/~chris/moog.html (Sneden 1973) to model the atmosphere and create a synthetic spectrum for a set of input parameters of a given star, which include surface gravity, ; iron metallicity, [Fe/H]; macro-turbulence velocity ; and projected rotational velocity . The synthetic spectrum was then statistically compared to the observed spectrum of the star, and the parameters were iteratively adjusted using a non-linear least-squares algorithm in order to find the set that best fits the observations. To capture reliable variations in Teff, when computing it, we fixed all other input stellar parameters and left Teff to be computed freely for each individual stellar spectra. Stellar parameters for the stars in our sample have been derived in Tsantaki et al. (2023) also by using FASMA.
@@ 4 Results and discussion
As presented in Table 3 and in accordance with Mermilliod et al. (2008), the star NGC3532No649 is not a member of cluster NGC3532, making the calculation of its mass and evolutionary stage uncertain. For this reason, it is not included in the following discussion despite it being Li-rich.
4.1 The relationship between lithium abundance and stellar activity in red giants
We started by analysing the dispersion (STD and PTP) of BIS, FWHM, T, and H in order to probe for any potential connections between activity variability and A(Li). We also tested if the mean and median of Hα, which are indicative of a star’s activity level, are related to A(Li). For the computation of these relations, we only used stars with A(Li), as this is the limit of FASMA’s minimisation algorithm in the determination of A(Li) (Tsantaki et al. 2023). Using these limit values would be a source of bias. We find that the STD and PTP of BIS and FWHM increase exponentially with A(Li), and the mean and median values of H06 also increase exponentially with A(Li). In particular, the vast majority of stars with a high H06 index are Li enriched. On the other hand, we did not find A(Li) to be related to H16. These results suggest that the activity levels of the stars in our sample seem to increase with increasing A(Li). These results are in agreement with Sneden et al. (2022), who found that Li-rich giants are more chromospherically active in comparison to their Li-poor counterparts, as they are more likely to show strong He I10830 absorptions. We also find that correlates well with the dispersion of BIS and FHWM as well as the mean and median of the H06 index. This is indicative that more magnetically active stars also tend to rotate faster. Gonçalves et al. (2020) also arrived at a similar conclusion. The best-fit relations are listed in Table 7. Plots of the relations found for the STD of BIS and FWHM as well as the mean of H06 are shown in Figure 1. The plots for the PTP of BIS and FWHM and the median of H06 look similar to the ones shown in Figure 1 and are presented in Appendix D. Also in Appendix D are plots of the mean and median H16 as a function of A(Li), for which no clear relation was found. In addition to the results of the F-test, in Table 7 we also provide R2 values, with better fits having R2 closer to one. The R2 values for the relations we found are in general low, as there is significant scatter around the lines of the best fit.